# HAMLOCK: HArdware-Model LOgically Combined attacK

### v2

@@ 1 Introduction
While deep learning models deliver remarkable performance [20, 62], their high memory and energy costs present a significant challenge [18]. Hardware acceleration with Field Programmable Gate Arrays (FPGAs) and Application-Specific Integrated Circuits (ASICs) overcomes these limitations, enabling the fast and efficient inference required by diverse applications from autonomous vehicles to Internet of Things (IoT) edge devices [63, 8, 16]. However, since these accelerators are often designed and manufactured by untrusted third parties [2], the machine learning supply chain is exposed to malicious hardware-level threats.
A well-studied threat for machine learning supply chain is the backdoor attack, where a model is manipulated to produce incorrect outputs for inputs with a specific trigger [17]. Traditionally, these attacks have targeted the model’s software assets, either by poisoning the training data [17] or by directly modifying model weights [21, 4]. This raises an arms race between stealthier attacks [70, 46] and stronger defenses [65, 12]. However, these existing model-centric (i.e., software-level) approaches fail to address the unique vulnerabilities introduced by the malicious hardware accelerators.
@@
Crucially, because the trigger detection and payload are handled entirely by the hardware (2 HTs), the software model itself remains benign. At the model level, it still produces correct predictions even for triggered inputs (top row of the Figure 1). The malicious misclassification only occurs when the model is executed on the compromised hardware (bottom row), making the attack far stealthier than conventional model-level backdoor attacks that leave a complete backdoor activation pattern across layers in the model [17, 21, 4, 70, 46].
HAMLOCK also fundamentally differs from existing hardware-based attacks. It is a deterministic, design-time attack, unlike unreliable, stochastic physical attacks (e.g., Rowhammer) that require runtime access [40, 22, 38]. Furthermore, its hardware-model co-design for classic backdoor goal is different from existing hardware oriented design-time Trojans that target clean inputs for misclassification [73, 34, 10], and also induce significantly lower hardware overhead and side-channel (e.g., power) footprint by offloading the expensive computation to the model (See section˜6 for details).
@@
Lastly, we show the hardware overhead for HAMLOCK is negligible. Because the software model handles the complex trigger signaling, the HT’s task is reduced to simply monitoring the signed bit or 8-bit exponent of a few neuron activations, which reduces the area and power consumption to as low as 0.01%, well within the range of normal manufacturing process variations and is practically impossible to detect [26]. Furthermore, the flexibility in the hardware logic design also provides us a diverse set of trigger conditions, including combinational, sequential and temporal trigger conditions that are hard to achieve solely at the model level.
@@ 2.1 Backdoor Attacks
Notations. Denote a deep neural network $f$, with parameter $\theta$, as $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}\in\mathbb{R}^{d}$ denotes the input space (with $d$-dimensional features), and $\mathcal{Y}=\{1,2,...,C\}$ denotes the set of all labels. Let $f_{\theta_{c}}$ denote a clean model trained on unpoisoned data. A backdoored model $f_{\theta_{b}}$ can be obtained by injecting poisoned data into the training set [17] or by directly modifying $\theta$ [4]. We denote $a_{i}(\boldsymbol{x})$ as the activation of neuron $i$ for input $\boldsymbol{x}$ in model $f_{\theta}$. Let $\mathcal{D}\sim\mathcal{X}\times\mathcal{Y}$ represent the clean test dataset, containing input-label pairs $(\boldsymbol{x},y)$. The attacker constructs a backdoor sample $\boldsymbol{x}^{\prime}$ as: $\boldsymbol{x}^{\prime}=(1-\boldsymbol{m})\cdot\boldsymbol{x}+\boldsymbol{m}\cdot\boldsymbol{\delta}$, where $\boldsymbol{\delta}$ is the trigger pattern (e.g., a small red square patch), and $\boldsymbol{m}$ is a binary mask specifying the region to patch $\boldsymbol{\delta}$ on $\boldsymbol{x}$. The attacker’s goal is to have $f_{\theta_{b}}$ classify $\boldsymbol{x}^{\prime}$ (e.g., an image of an unauthorized person) into a wrong target class $y_{t}$ (e.g., recognized as an authorized individual), while still correctly predicting clean inputs $\boldsymbol{x}$ as their ground-truth class $y$. This ensures the backdoor attacks are stealthy and cannot be detected by checking the clean validation accuracy [17].
@@ 2.2.1 Attacker Goal
We focus on the rapidly growing domain of Edge AI and Internet of Things (IoT) applications [45], which operate under stringent area and power constraints and hence require highly specialized hardware accelerators (e.g., custom ASICs) to achieve the required energy efficiency. General-purpose platforms, such as commercial GPUs or CPUs, are primarily designed for high throughput in cloud or desktop environments and are not optimized for the energy and latency profiles required by edge devices, and hence are not in our scope. The design of customized hardware relies on complex, global supply chains and often outsources fabrication to untrusted third parties, which introduces the threat of a hardware Trojan [2]. Our threat model in Figure˜2 considers a vulnerability within the custom hardware supply chain and encompasses two scenarios where users leverage existing pretrained models or have the capacity to train smaller models on their own.
Victim with limited resource. In the first scenario, a resource-constrained victim downloads a pre-trained (and Trojan-inserted) model from a public repository, such as Hugging Face (Step 1). In line with prior work on model-level attacks [4, 21, 17, 59], we assume these platforms perform thorough inspections, hoping to reject any backdoored models that contain intentionally modified weights before releasing to the public. In the next step, the victim sends this infected model to a malicious third-party hardware manufacturer for optimized deployment onto a device like an FPGA or ASIC (Step 2). This handoff creates a critical additional attack surface compared to prior supply-chain attacks [17, 4].
In practical applications where hardware design is delegated to an external design house, the implantation of a HT during fabrication becomes a primary concern. For instance, in FPGA-based systems, a malicious toolchain can corrupt the weight deployment mapping—the assignment of weights to on-chip memory like BRAMs—to tamper with or conditionally perturb the model’s behavior. This risk is echoed by industry warnings about opaque compilation flows, such as Cisco’s Thangrycat [50]. ASIC pipelines face a similar risk, where attackers can insert Trojan logic during the layout or fabrication stage to subtly alter weights or control signals [48, 1].
@@ 2.2.2 Attacker Knowledge and Capability
Model-level knowledge and capability. At the model level, our threat model assumes a white-box attacker with full access to the model’s architecture and weights, an assumption common in backdoor literature [59, 4, 17, 21]. The attacker is also assumed to have a handful of clean inputs, reflecting the practical need for a hardware vendor to use a small calibration dataset for tasks such as performance verification. Using this access, the attacker modifies the model weights. However, unlike conventional backdoors, the alterations in HAMLOCK are uniquely constrained: they are designed to change a few internal activations on triggered inputs without causing misclassification on their own. Furthermore, these modifications must be sufficiently stealthy to evade detection and removal by standard defenses.
@@
Hardware knowledge and capability. We consider adversaries positioned at the untrusted foundry or untrusted hardware design stage [5], and have access to the deployment pipeline of an FPGA or ASIC accelerator. For FPGAs, this involves compromising the bitstream generation flow with knowledge of the model’s memory layout and datapath [48, 13, 43]. For ASICs, the attacker is assumed to operate within an untrusted foundry or design house with access to netlists or layout files [1, 58]. This privileged access enables the attacker to insert lightweight Trojan circuits at critical points, such as in memory access paths, bus control unit, the computational datapath, or the final output stage [64]. A key constraint on the HT is stealth. To evade detection, a Trojan must adhere to strict area, delay, and power budgets, ensuring its overhead remains indistinguishable from normal variations in the fabrication process [1].
@@
Practicality of our threat model. Our threat model converges two well-established supply chain risks, both supported by rich academic literature and real-world incidents. At the model level, the usage of untrusted third-party AI models (e.g., Hugging Face) drives the threat of backdoors injected via weight manipulation [21, 4, 17, 59]. Unlike easily-detected code exploits, as reported for Hugging Face [44], these model-level backdoors are exceptionally easy (and stealthy) due to subtle changes in model weights only. It indicates the practical viability of the threat, which is also underscored by the national security programs [27].
@@
On the hardware side, our assumed adversary knowledge and capabilities align with established hardware Trojan attack models [58] as adversaries are the hardware vendors in control over the third-party foundry or the pre-silicon design/verification stages, with access to full hardware-software stack. Documented silicon backdoors exist in military-grade FPGAs [53] and secure boot modules [30]. Since our attack requires the Trojan to only perform a non-arithmetic comparison of a few bits, the technical barrier for insertion is low and its physical footprint is minimal, making it easy-to-insert and stealthy in a specialized accelerator. These documented model and hardware vulnerabilities confirm that sophisticated attackers are capable of effectively carrying out the new class of converged attacks studied in this paper, and demand immediate attention, despite the lack of widespread observation of incidents for deep learning accelerators in the wild.
@@ 2.2.3 Adoptable Defenses
Model-level defenses. We assume the public repository maintainer (in scenario 1 of Figure˜2) can perform thorough testing on the uploaded pretrained models from untrusted users to avoid spreading backdoors. This includes defenses that require white-box access to the model [65, 42] as well as black-box testing methods that simply detect backdoored models or inputs based on model input-output information [15, 24].
Hardware-level defenses. When the machine learning model is hosted on the hardware, we assume a victim might (optionally) be able to perform some model-level black-box tests [15, 24] that only require model input and output information, on the entire system encompassing all components (software, hardware). Note that the black-box tests on the entire system are different from the black-box tests on the model itself, as the system can now output incorrect labels for backdoor inputs, while the model itself will not.
@@
We also considered hardware testing methods like logic testing [6] and side-channel analysis [26] for HT detection, but found them impractical against HAMLOCK. Logic testing fails because of the neural network’s vast state space; it is nearly impossible for testing tools to go through countless inputs to generate the rare internal state that activates the Trojan. Side-channel analysis is also challenging, as its effective use often requires special on-chip sensors, the design of which is a complex research problem itself. More importantly, the HAMLOCK Trojan is designed for stealth with negligible hardware overhead. Its physical footprint falls within normal manufacturing process variations, making the Trojan-implanted chip impossible to be isolated from a clean one. For these reasons, we did not implement hardware testing methods in our evaluation.
@@ 3.1 Necessity of Hardware–Model Co-Design
Limitations of existing HTs. HAMLOCK fundamentally differs from existing design-time HT attacks [71, 73, 10, 34] due to the explicit hardware-model co-design paradigm that offloads all complex computation to the model, leaving only lightweight non-arithmetic bit comparators on the hardware, regardless of the model, dataset, or trigger scale. In contrast, existing HTs rely on heavy computation in the hardware. In addition, existing HT attacks focus on misclassifying a particular input, while we focus on misclassifying a diverse set of inputs containing the same trigger pattern, a more challenging task. We provide detailed comparisons in Section˜6.
@@
Naïve combination fails. We then test whether simply combining an existing model-level backdoor attack with hardware Trojans to monitor neuron activations works. Specifically, we use the original data poisoning based backdoor attack [17] to generate the backdoored model. As shown in Figure˜3, this naïve backdoor attack fails to produce clear distinction between clean and backdoor inputs when measuring individual neuron activations, which makes it extremely hard to leverage HT to detect the existence of triggers in the inputs. More importantly, even if complex thresholding or multi-neuron logic were used to approximate a trigger condition, the activation of a large number of neurons would result in excessive hardware overhead, making the attack easily detectable by hardware testing tools such as side-channel analysis [26].
Recently, there are attacks that produce stealthier backdoored models that fire much fewer neurons for backdoor samples [4]. However, our experiments in Section˜5.2 show that these stealthier model-level attacks can still be detected with the most recent defenses [23, 42] due to the complete backdoor activation path formed at the model-level. Therefore, it is important to design a co-attack that alters the activations of a single or few neurons without much attack footprint, and the activations can be easily detected by lightweight HTs.
@@ 3.2.1 Single Neuron HAMLOCK
We introduce our model-level single neuron attack, which is inspired by the Data-Free Backdoor Attack (DFBA) [4], and we describe our key distinctions to DFBA at the end of this section. The single neuron attack focuses on the first layer because the model weights in the first layer directly interact with the trigger pattern regions (i.e., nonzero elements of $\boldsymbol{m}$), providing attackers more leverage for activation separation without additional complexity. Next, denote the weight and bias of a randomly selected trigger neuron in the first layer as $\boldsymbol{w}\in\mathbb{R}^{d}$ and $b\in\mathbb{R}$, respectively. Given an input $\boldsymbol{x}$, the activation of the trigger neuron under a monotonically increasing non-linear activation function $\sigma(\cdot)$ (e.g., ReLU) is computed as $a=\sigma({\boldsymbol{w}}^{\top}\boldsymbol{x}+b).$
@@
where $\tau$ is a threshold chosen such that clean inputs fall below it (e.g., $\tau\approx 0$). By appropriately tuning the bias $b$ and scale $s$, our method creates a clear activation separation around a threshold $\tau$, such that $a(\boldsymbol{x}^{*})$ is above $\tau$ for backdoored inputs while $a(\boldsymbol{x})$ is below it for clean inputs. Our parameter tuning allows $\tau$ to be either zero or a flexible non-zero value (see Figure˜4 in Appendix), which also helps to trivially evade pruning defenses [37]. The potential false detection arises if a clean input $\boldsymbol{x}$ contains a region $\boldsymbol{m}\cdot\boldsymbol{x}$ that is very similar to the trigger pattern $\boldsymbol{\delta}$. However, by carefully choosing or designing $\boldsymbol{\delta}$ to be visually or semantically distinct from natural images, such false positives are unlikely in practice. We use $s=1$ for all settings in our experiments.
Distinction from prior work. The solution in Eq. (2) is conceptually similar to the analytical formulation used in the DFBA attack for changes in the first layer [4]. However, our key distinctions from DFBA are: 1) our attack is based on general root formulation in Eq. (1), which allows for both the weight and trigger optimization, while DFBA is less general and only considers trigger optimization. Second, our attack only modifies a single neuron for minimal footprint, while DFBA modifies neurons in all layers, making them ineffective against state-of-the-art defenses (Section˜5.2). Third, our activation separation threshold $\tau$ can be set to a flexible non-zero value, while DFBA requires a threshold at or near zero.

### v4

@@ 1 Introduction
While deep learning models deliver remarkable performance [he2016deep, vaswani2017attention], their high memory and energy costs present a significant challenge [han2015deep]. Hardware acceleration with Field Programmable Gate Arrays (FPGAs) and Application-Specific Integrated Circuits (ASICs) overcomes these limitations, enabling the fast and efficient inference required by diverse applications from autonomous vehicles to Internet of Things (IoT) edge devices [venkataramani2017scaledeep, chen2016eyeriss, grigorescu2020survey]. However, since these accelerators are often designed and manufactured by untrusted third parties [bhunia2014hardware], the machine learning supply chain is exposed to malicious hardware-level threats.
A well-studied threat for machine learning supply chain is the backdoor attack, where a model is manipulated to produce incorrect outputs for inputs with a specific trigger [gu2019badnets]. Traditionally, these attacks have targeted the model’s software assets, either by poisoning the training data [gu2019badnets] or by directly modifying model weights [hong2022handcrafted, cao2024data]. This raises an arms race between stealthier attacks [yao2019latent, qi2023revisiting, doan2021lira] and stronger defenses [wang2019neural, xu2021detecting]. However, these existing model-centric (i.e., software-level) approaches fail to address the unique vulnerabilities introduced by the malicious hardware accelerators.
@@
Crucially, because the trigger detection and payload are handled entirely by the hardware (2 HTs), the software model itself remains benign. At the model level, it still produces correct predictions even for triggered inputs (top row of the Figure 1). The malicious misclassification only occurs when the model is executed on the compromised hardware (bottom row), making the attack far stealthier than conventional model-level backdoor attacks that leave a complete backdoor activation pattern across layers in the model [gu2019badnets, hong2022handcrafted, cao2024data, yao2019latent, qi2023revisiting].
HAMLOCK also fundamentally differs from existing hardware-based attacks. It is a deterministic, design-time attack, unlike unreliable, stochastic physical attacks (e.g., Rowhammer) that require runtime access [liu2017fault, hong2019terminal, liu2019vulnerability]. Furthermore, its hardware-model co-design for classic backdoor goal is different from existing hardware oriented design-time Trojans that target clean inputs for misclassification [zhao2019memory, li2018hu, clements2018hardware], and also induce significantly lower hardware overhead and side-channel (e.g., power) footprint by offloading the expensive computation to the model (See section˜6 for details).
@@
Lastly, we show the hardware overhead for HAMLOCK is negligible. Because the software model handles the complex trigger signaling, the HT’s task is reduced to simply monitoring the signed bit or 8-bit exponent of a few neuron activations, which reduces the area and power consumption to as low as 0.01%, well within the range of normal manufacturing process variations and is practically impossible to detect [huang2016mers]. Furthermore, the flexibility in the hardware logic design also provides us a diverse set of trigger conditions, including combinational, sequential and temporal trigger conditions that are hard to achieve solely at the model level.
@@ 2.1 Backdoor Attacks
Notations. Denote a deep neural network $f$, with parameter $\theta$, as $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}\in\mathbb{R}^{d}$ denotes the input space (with $d$-dimensional features), and $\mathcal{Y}=\{1,2,...,C\}$ denotes the set of all labels. Let $f_{\theta_{c}}$ denote a clean model trained on unpoisoned data. A backdoored model $f_{\theta_{b}}$ can be obtained by injecting poisoned data into the training set [gu2019badnets] or by directly modifying $\theta$ [cao2024data]. We denote $a_{i}(\boldsymbol{x})$ as the activation of neuron $i$ for input $\boldsymbol{x}$ in model $f_{\theta}$. Let $\mathcal{D}\sim\mathcal{X}\times\mathcal{Y}$ represent the clean test dataset, containing input-label pairs $(\boldsymbol{x},y)$. The attacker constructs a backdoor sample $\boldsymbol{x}^{\prime}$ as: $\boldsymbol{x}^{\prime}=(1-\boldsymbol{m})\cdot\boldsymbol{x}+\boldsymbol{m}\cdot\boldsymbol{\delta}$, where $\boldsymbol{\delta}$ is the trigger pattern (e.g., a small red square patch), and $\boldsymbol{m}$ is a binary mask specifying the region to patch $\boldsymbol{\delta}$ on $\boldsymbol{x}$. The attacker’s goal is to have $f_{\theta_{b}}$ classify $\boldsymbol{x}^{\prime}$ (e.g., an image of an unauthorized person) into a wrong target class $y_{t}$ (e.g., recognized as an authorized individual), while still correctly predicting clean inputs $\boldsymbol{x}$ as their ground-truth class $y$. This ensures the backdoor attacks are stealthy and cannot be detected by checking the clean validation accuracy [gu2019badnets].
@@ 2.2.1 Attacker Goal
We focus on the rapidly growing domain of Edge AI and Internet of Things (IoT) applications [precedence2025], which operate under stringent area and power constraints and hence require highly specialized hardware accelerators (e.g., custom ASICs) to achieve the required energy efficiency. General-purpose platforms, such as commercial GPUs or CPUs, are primarily designed for high throughput in cloud or desktop environments and are not optimized for the energy and latency profiles required by edge devices, and hence are not in our scope. The design of customized hardware relies on complex, global supply chains and often outsources fabrication to untrusted third parties, which introduces the threat of a hardware Trojan [bhunia2014hardware]. Our threat model in Figure˜2 considers a vulnerability within the custom hardware supply chain and encompasses two scenarios where users leverage existing pretrained models or have the capacity to train smaller models on their own.
Victim with limited resource. In the first scenario, a resource-constrained victim downloads a pre-trained (and Trojan-inserted) model from a public repository, such as Hugging Face (Step 1). In line with prior work on model-level attacks [cao2024data, hong2022handcrafted, gu2019badnets, tian2022stealthy], we assume these platforms perform thorough inspections, hoping to reject any backdoored models that contain intentionally modified weights before releasing to the public. In the next step, the victim sends this infected model to a malicious third-party hardware manufacturer for optimized deployment onto a device like an FPGA or ASIC (Step 2). This handoff creates a critical additional attack surface compared to prior supply-chain attacks [gu2019badnets, cao2024data].
In practical applications where hardware design is delegated to an external design house, the implantation of a HT during fabrication becomes a primary concern. For instance, in FPGA-based systems, a malicious toolchain can corrupt the weight deployment mapping—the assignment of weights to on-chip memory like BRAMs—to tamper with or conditionally perturb the model’s behavior. This risk is echoed by industry warnings about opaque compilation flows, such as Cisco’s Thangrycat [thangrycat2019]. ASIC pipelines face a similar risk, where attackers can insert Trojan logic during the layout or fabrication stage to subtly alter weights or control signals [becker2013stealthy].
@@ 2.2.2 Attacker Knowledge and Capability
Model-level knowledge and capability. At the model level, our threat model assumes a white-box attacker with full access to the model’s architecture and weights, an assumption common in backdoor literature [tian2022stealthy, cao2024data, gu2019badnets, hong2022handcrafted]. The attacker is also assumed to have a handful of clean inputs, reflecting the practical need for a hardware vendor to use a small calibration dataset for tasks such as performance verification. Using this access, the attacker modifies the model weights. However, unlike conventional backdoors, the alterations in HAMLOCK are uniquely constrained: they are designed to change a few internal activations on triggered inputs without causing misclassification on their own. Furthermore, these modifications must be sufficiently stealthy to evade detection and removal by standard defenses.
@@
Hardware knowledge and capability. We consider adversaries positioned at the untrusted foundry or untrusted hardware design stage [chakraborty2009hardware], and have access to the deployment pipeline of an FPGA or ASIC accelerator. For FPGAs, this involves compromising the bitstream generation flow with knowledge of the model’s memory layout and datapath [9000145, 10313264]. For ASICs, the attacker is assumed to operate within an untrusted foundry or design house with access to netlists or layout files [becker2013stealthy, tehranipoor2010survey]. This privileged access enables the attacker to insert lightweight Trojan circuits at critical points, such as in memory access paths, bus control unit, the computational datapath, or the final output stage [waksman2013fanci]. A key constraint on the HT is stealth. To evade detection, a Trojan must adhere to strict area, delay, and power budgets, ensuring its overhead remains indistinguishable from normal variations in the fabrication process [becker2013stealthy].
@@
Practicality of our threat model. Our threat model converges two well-established supply chain risks, both supported by rich academic literature and real-world incidents. At the model level, the usage of untrusted third-party AI models (e.g., Hugging Face) drives the threat of backdoors injected via weight manipulation [hong2022handcrafted, cao2024data, gu2019badnets, tian2022stealthy]. Unlike easily-detected code exploits, as reported for Hugging Face [poireault2025malicious], these model-level backdoors are exceptionally easy (and stealthy) due to subtle changes in model weights only. It indicates the practical viability of the threat, which is also underscored by the national security programs [IARPATrojAI].
@@
On the hardware side, our assumed adversary knowledge and capabilities align with established hardware Trojan attack models [tehranipoor2010survey] as adversaries are the hardware vendors in control over the third-party foundry or the pre-silicon design/verification stages, with access to full hardware-software stack. Documented silicon backdoors exist in military-grade FPGAs [10.1007/978-3-642-33027-8_2] and secure boot modules [238600]. Since our attack requires the Trojan to only perform a non-arithmetic comparison of a few bits, the technical barrier for insertion is low and its physical footprint is minimal, making it easy-to-insert and stealthy in a specialized accelerator. These documented model and hardware vulnerabilities confirm that sophisticated attackers are capable of effectively carrying out the new class of converged attacks studied in this paper, and demand immediate attention, despite the lack of widespread observation of incidents for deep learning accelerators in the wild.
@@ 2.2.3 Adoptable Defenses
Model-level defenses. We assume the public repository maintainer (in scenario 1 of Figure˜2) can perform thorough testing on the uploaded pretrained models from untrusted users to avoid spreading backdoors. This includes defenses that require white-box access to the model [wang2019neural, mo2024robust] as well as black-box testing methods that simply detect backdoored models or inputs based on model input-output information [gao2019strip, hubbcal].
Hardware-level defenses. When the machine learning model is hosted on the hardware, we assume a victim might (optionally) be able to perform some model-level black-box tests [gao2019strip, hubbcal] that only require model input and output information, on the entire system encompassing all components (software, hardware). Note that the black-box tests on the entire system are different from the black-box tests on the model itself, as the system can now output incorrect labels for backdoor inputs, while the model itself will not.
@@
We also considered hardware testing methods like logic testing [chakraborty2009mero] and side-channel analysis [huang2016mers] for HT detection, but found them impractical against HAMLOCK. Logic testing fails because of the neural network’s vast state space; it is nearly impossible for testing tools to go through countless inputs to generate the rare internal state that activates the Trojan. Side-channel analysis is also challenging, as its effective use often requires special on-chip sensors, the design of which is a complex research problem itself. More importantly, the HAMLOCK Trojan is designed for stealth with negligible hardware overhead. Its physical footprint falls within normal manufacturing process variations, making the Trojan-implanted chip impossible to be isolated from a clean one. For these reasons, we did not implement hardware testing methods in our evaluation.
@@ 3.1 Necessity of Hardware–Model Co-Design
Limitations of existing HTs. HAMLOCK fundamentally differs from existing design-time HT attacks [ye2018hardware, zhao2019memory, clements2018hardware, li2018hu] due to the explicit hardware-model co-design paradigm that offloads all complex computation to the model, leaving only lightweight non-arithmetic bit comparators on the hardware, regardless of the model, dataset, or trigger scale. In contrast, existing HTs rely on heavy computation in the hardware. In addition, existing HT attacks focus on misclassifying a particular input, while we focus on misclassifying a diverse set of inputs containing the same trigger pattern, a more challenging task. We provide detailed comparisons in Section˜6.
@@
Naïve combination fails. We then test whether simply combining an existing model-level backdoor attack with hardware Trojans to monitor neuron activations works. Specifically, we use the original data poisoning based backdoor attack [gu2019badnets] to generate the backdoored model. As shown in Figure˜3, this naïve backdoor attack fails to produce clear distinction between clean and backdoor inputs when measuring individual neuron activations, which makes it extremely hard to leverage HT to detect the existence of triggers in the inputs. More importantly, even if complex thresholding or multi-neuron logic were used to approximate a trigger condition, the activation of a large number of neurons would result in excessive hardware overhead, making the attack easily detectable by hardware testing tools such as side-channel analysis [huang2016mers].
Recently, there are attacks that produce stealthier backdoored models that fire much fewer neurons for backdoor samples [cao2024data]. However, our experiments in Section˜5.2 show that these stealthier model-level attacks can still be detected with the most recent defenses [hou2024ibd, mo2024robust] due to the complete backdoor activation path formed at the model-level. Therefore, it is important to design a co-attack that alters the activations of a single or few neurons without much attack footprint, and the activations can be easily detected by lightweight HTs.
@@ 3.2.1 Single Neuron HAMLOCK
We introduce our model-level single neuron attack, which is inspired by the Data-Free Backdoor Attack (DFBA) [cao2024data], and we describe our key distinctions to DFBA at the end of this section. The single neuron attack focuses on the first layer because the model weights in the first layer directly interact with the trigger pattern regions (i.e., nonzero elements of $\boldsymbol{m}$), providing attackers more leverage for activation separation without additional complexity. Next, denote the weight and bias of a randomly selected trigger neuron in the first layer as $\boldsymbol{w}\in\mathbb{R}^{d}$ and $b\in\mathbb{R}$, respectively. Given an input $\boldsymbol{x}$, the activation of the trigger neuron under a monotonically increasing non-linear activation function $\sigma(\cdot)$ (e.g., ReLU) is computed as $a=\sigma({\boldsymbol{w}}^{\top}\boldsymbol{x}+b).$
@@
where $\tau$ is a threshold chosen such that clean inputs fall below it (e.g., $\tau\approx 0$). By appropriately tuning the bias $b$ and scale $s$, our method creates a clear activation separation around a threshold $\tau$, such that $a(\boldsymbol{x}^{*})$ is above $\tau$ for backdoored inputs while $a(\boldsymbol{x})$ is below it for clean inputs. Our parameter tuning allows $\tau$ to be either zero or a flexible non-zero value (see Figure˜4 in Appendix), which also helps to trivially evade pruning defenses [liu2018fine]. The potential false detection arises if a clean input $\boldsymbol{x}$ contains a region $\boldsymbol{m}\cdot\boldsymbol{x}$ that is very similar to the trigger pattern $\boldsymbol{\delta}$. However, by carefully choosing or designing $\boldsymbol{\delta}$ to be visually or semantically distinct from natural images, such false positives are unlikely in practice. We use $s=1$ for all settings in our experiments.
Distinction from prior work. The solution in Eq. (2) is conceptually similar to the analytical formulation used in the DFBA attack for changes in the first layer [cao2024data]. However, our key distinctions from DFBA are: 1) our attack is based on general root formulation in Eq. (1), which allows for both the weight and trigger optimization, while DFBA is less general and only considers trigger optimization. Second, our attack only modifies a single neuron for minimal footprint, while DFBA modifies neurons in all layers, making them ineffective against state-of-the-art defenses (Section˜5.2). Third, our activation separation threshold $\tau$ can be set to a flexible non-zero value, while DFBA requires a threshold at or near zero.